# NPHardEval: Dynamic Benchmark on Reasoning Ability of Large Language Models via Complexity Classes

### 25 Dec 2023 v2

@@ Abstract
Complex reasoning ability is one of the most important features of current LLMs, which has also been leveraged to play an integral role in complex decision-making tasks. Therefore, the investigation into the reasoning capabilities of Large Language Models (LLMs) is critical: numerous benchmarks have been established to assess the reasoning abilities of LLMs. However, current benchmarks are inadequate in offering a rigorous evaluation of the full extent of reasoning abilities that LLMs are capable of achieving. They are also prone to the risk of overfitting, as these benchmarks, being publicly accessible and static, allow models to potentially tailor their responses to specific benchmark metrics, thereby inflating their performance. Addressing these limitations, our research introduces a new benchmark, named NPHardEval. This benchmark is designed to evaluate the reasoning abilities of LLMs across a broad spectrum of 900 algorithmic questions, extending up to the NP-Hard complexity class. These questions are meticulously chosen to represent a wide range of complexity class below the NP-hard complexity class, offering a rigorous measure of the reasoning ability of LLMs. Through this study, we shed light on the current state of reasoning in LLMs, providing an objective and rigorous perspective through the comparison of LLMs’ performance across complex classes. Our findings contribute significantly to understanding the current capabilities of LLMs in reasoning tasks and lay the groundwork for future advancements in enhancing the reasoning abilities of these models. Moreover, this benchmark is designed with a dynamic update mechanism, where the datapoints are refreshed on a monthly basis. Such regular updates play a crucial role in mitigating the risk of LLMs overfitting to the benchmark, promoting a more accurate and reliable assessment of their reasoning capabilities. The benchmark dataset and code of NPHardEval are available at https://github.com/casmlab/NPHardEval.
@@ 1 Introduction
    Reasoning Ability Gap among Foundation Models: How far away are current open-sourced models compared with closed-sourced models? We aim to quantitatively analyze the differences in performance between closed-source models (including as GPT 4 Turbo, Claude 2, GPT 3.5 Turbo, Claude Instant, and PaLM 2) and open-source models (including Vicuna-13b, Yi-34b, Mistra-7b, MPT-30b, and Phi-1.5). This comparison will shed light on the relative strengths and weaknesses of these models in complex reasoning scenarios.
@@ 2.1 Reasoning ability of LLMs
LLMs brown2020language ; chowdhery2023palm ; chung2022scaling have made significant advancements in natural language processing and related fields. Recent research underscores the unprecedented reasoning abilities of LLMshuang2022towards . It has been discussed that these models exhibit “emergent” behaviors, including the ability to “reason” when they are large enough wei2022emergent ; schaeffer2023emergent . By providing the models with the chain of thoughts with a simple prompt “Let’s think step by step”, these models are able to answer questions with explicit reasoning steps wei2022chain . This has sparked considerable interest in the community since reasoning ability is a hallmark of human intelligence. Various variations of chain-of-thought have been developed to prompt models’ reasoning abilitykojima2022large ; wang2022iteratively , such as tree of thoughtyao2023tree , graph of thoughtbesta2023graph , self-inspring techniquewang2023recmind .
@@ 2.2 Benchmarks of LLMs’ Performance
The advancement of LLMs has catalyzed the evolution of a range of general-purpose AI technologies, underscoring the importance of accurately assessing these models’ reasoning capabilities. Existing evaluation approaches predominantly rely on datasets comprising human-generated questions and their standard answers. For instance, MMLU hendrycks2020measuring and GAOKAO zhang2023evaluating both utilize human exam questions in their automated evaluations. Additionally, datasets such as the French National Math Exam, Hungarian National High School Exam [^1] https://huggingface.co/datasets/keirp/hungarian_national_hs_finals_exam, and GHOST (Graduate-Level High-Order Skill Tests)frieder2023mathematical are utilized to assess LLMs’ reasoning proficiency. These sources aim to ensure the absence of data leakage. Nonetheless, the requirement for manual verification of answers in these datasets limits their practical utility. In general, these datasets are commonly employed as benchmarks in the field; however, they lack a quantitative metric for assessing the difficulty level of the questions and the extent of reasoning necessary to answer them. This absence of precise measurement criteria results in a limited understanding of the logical reasoning capabilities of large language models.
@@ 3.2 Difficulty Level for Tasks
NPHardEval categorizes the challenges it presents into a hierarchy of difficulty, spanning from the simplest to the most complex. This structure is divided into 10 distinct levels of difficulty for each task, with the initial level being designed as the most basic challenge that a Large Language Model (LLM) might face. This gradation allows for a nuanced assessment of an LLM’s problem-solving abilities across a spectrum of increasingly complex tasks. For instance, the GCP-D problem has difficulty levels 1 to 10 with questions of 6, 8, 10, 12, 14, 16, 18, 20, 22, and 24 average edges and 6, 7, 8, 9, 10, 11, 12, 13, 14, and 15 nodes. Beginning with graphs of 6 nodes and 6 edges, each subsequent level incorporates an additional 2 edges and 1 node, culminating in graphs of 24 edges and 15 nodes at the most challenging level.
@@ 4 Experimental Setting
This section presents the experiment setting to answer the three research questions. Our approach assessed 10 distinct LLMs, with a dichotomy between five proprietary (closed-source) models and five open-source models, including GPT 4 Turbo, Claude 2, GPT 3.5 Turbo, Claude Instant 1.2, PaLM 2, Vicuna-13b, Yi-34b, Mistra-7b, MPT-30b, and Phi-1.5.
@@ 4.2 Experiment 2: Model Performance across Task Complexity
4.3 Experiment 3: Comparative Analysis of Learnability by In-context Learning
@@ 5.1 Reasoning Ability of Foundation Models
Experiment 1 focuses on a comprehensive comparison among various foundation models. In Figure 2, we present the overall zero-shot accuracy for each problem, providing a visual representation of the performance of different models. Our observations reveal that closed-source models generally demonstrate higher accuracy and a lower rate of failure compared to their open-source counterparts. Notably, GPT-4 often emerges as the frontrunner in performance across the majority of tasks, indicating its superior problem-solving capabilities in zero-shot settings. Within the realm of open-source models, Yi-34b and Mistral-7b distinguish themselves by significantly outperforming other models in this category. This disparity highlights a notable performance gap and suggests that these two models possess more advanced reasoning abilities compared to other open-source options.
@@
As Figure 3(a) shows, upon analysis of the weighted accuracy for different models across problem complexities, we observed a general trend where all models experienced a decrease in accuracy as problem complexity increased. Notably, there are two detailed findings for overall reasoning ability change. First, regarding the performance decay speed, among the 10 models we tested, the average performance demonstrated a higher accuracy at the P and NP-Complete complexity levels (both with a weighted accuracy of 0.26) but saw a sharper decline as the problems became more complex when proceed to the NP-hard level (with a weighted accuracy of 0.03). There is a performance decay on average when models are test against NP-Hard problems. Second, close-source models usually perform better than open source models – there are more triangles in the upper locations than than squares in Figure 3(a).
Figure 3: Model performance on different complexity problems: (a) weighted accuracy (b) (weighted) failure rate. Open models are denoted in squares and close models are denoted in triangles.
@@ 5.2 Reasoning Performance across Task Complexity and Difficulty Levels
Figure 4: Model’s performance on each complexity level. (a) GPT 4 Turbo. (b) Claude 2. (c) GPT 3.5 Turbo. (d) Claude Instant 1.2. (e) PaLM 2. (f) Yi-34b. (g) Mistra-7b. (h) MPT-30b. (i) Vicuna-13b. (j) Phi-1.5.
Figure 5 presents the accuracy of each model across various problems associated with P, NP-Complete, and NP-Hard complexities. Regarding P complexity, notable differences emerged among the models. GPT 3.5 Turbo, GPT 4 Turbo, and Yi-34b models exhibited significantly superior performance on the SAS problem compared to the other two problems. GPT 3.5 Turbo, Yi-34b, and Vicuna-13b models demonstrated markedly better performance on the EDP problem compared to the SPP problem. Only the Vicuna-13b model displayed slightly better performance, although not significant, on the EDP problem compared to SAS across all investigated models.
Other observations include: GPT 4 Turbo showcased very similar performance between the EDP and SPP problems, while Claude Instant 1.2 exhibited similar performance for all these three problems. Yi-34b, GPT 3.5 Turbo, and GPT 4 Turbo displayed remarkably high accuracy specifically for the SAS task. MPT-30b and Phi-1.5 showed very limited performance in identifying these three problems.
Regarding NP-Complete complexity, there are several observations to highlight. Still, neither MPT-30b nor Phi-1.5 could deliver any identification for problems in the NP-Complete complexity. In the case of GCP-D and TSP-D problems, the performance of these models varied significantly. Vicuna-13b and GPT 4 Turbo outperformed in the GCP-D problem compared to tsp-d, whereas Claude Instant 1.2, Claude 2, and PaLM 2 exhibited significantly better performance in TSP-D over GCP-D. On the other hand, models like Mistral-7b, Yi-34b, and GPT 3.5 Turbo showcased relatively similar performance between these two tasks. For the KSP task, only GPT 4 Turbo demonstrated promising performance, while the remaining models faltered.
@@
Figure 5: Model’s performance on tasks across complexity levels. (a) GPT 4 Turbo. (b) Claude 2. (c) GPT 3.5 Turbo. (d) Claude Instant 1.2. (e) PaLM 2. (f) Yi-34b. (g) Mistra-7b. (h) MPT-30b. (i) Vicuna-13b. (j) Phi-1.5.
@@ 5.3 Effects of Few-shot Examples’ Difficulty on Reasoning Ability Enhancement
In Experiment 3, we focused on the tasks of SAS and EDP to investigate the nature of the in-context learning capabilities of LLMs. This experiment empirically distinguishes between “learning” and “mimicking” as exhibited by LLMs during in-context learning scenarios. Our findings also revealed a clear dichotomy in the approach to learning and generalization from examples between closed-source and open-source models.
@@
Accuracy GPT 4 Turbo Claude 2 GPT 3.5 Turbo Claude Instant PaLM 2 Yi-34b Mistra-7b MPT-30b Vicuna-13b Phi-1.5
@@
Zeroshot 1.000 0.445 0.942 0.442 0.416 0.620 0.149 0.000 0.113 0.000
Fewshot (-5) 0.978 0.685 0.920 0.735 0.603 0.065 0.043 0.165 0.085 0.155
Fewshot (-4) 1.000 0.662 0.902 0.667 0.516 0.093 0.129 0.156 0.084 0.118
Fewshot (-3) 0.982 0.694 0.831 0.769 0.496 0.143 0.153 0.131 0.067 0.084
Fewshot (-2) 1.000 0.771 0.910 0.710 0.617 0.102 0.094 0.117 0.037 0.087
Fewshot (-1) 0.987 0.770 0.896 0.589 0.607 0.048 0.085 0.157 0.087 0.057
Fewshot (0) 0.984 0.671 0.846 0.651 0.660 0.598 0.413 0.258 0.089 0.098
Fewshot (1) 0.991 0.580 0.878 0.696 0.455 0.593 0.386 0.233 0.055 0.109
Fewshot (2) 1.000 0.675 0.829 0.587 0.656 0.647 0.296 0.175 0.067 0.056
Fewshot (3) 0.993 0.736 0.800 0.598 0.489 0.662 0.318 0.144 0.093 0.098
Fewshot (4) 1.000 0.729 0.869 0.580 0.471 0.638 0.287 0.269 0.053 0.106
Fewshot (5) 1.000 0.671 0.844 0.602 0.607 0.167 0.356 0.202 0.055 0.064
@@
Zeroshot 0.536 0.120 0.318 0.176 0.033 0.165 0.058 0.002 0.147 0.000
Fewshot (-5) 0.387 0.075 0.417 0.048 0.170 0.000 0.015 0.000 0.000 0.205
Fewshot (-4) 0.556 0.209 0.367 0.102 0.207 0.000 0.000 0.000 0.044 0.284
Fewshot (-3) 0.500 0.178 0.386 0.167 0.235 0.029 0.029 0.000 0.108 0.331
Fewshot (-2) 0.462 0.173 0.479 0.210 0.208 0.146 0.090 0.000 0.154 0.335
Fewshot (-1) 0.485 0.200 0.513 0.246 0.289 0.135 0.098 0.011 0.248 0.328
Fewshot (0) 0.518 0.209 0.564 0.253 0.238 0.282 0.182 0.022 0.164 0.293
Fewshot (1) 0.535 0.184 0.535 0.355 0.205 0.089 0.089 0.013 0.160 0.115
Fewshot (2) 0.545 0.209 0.544 0.238 0.196 0.195 0.042 0.015 0.093 0.087
Fewshot (3) 0.536 0.189 0.449 0.315 0.182 0.127 0.067 0.007 0.191 0.051
Fewshot (4) 0.538 0.209 0.507 0.305 0.200 0.247 0.095 0.000 0.129 0.000
Fewshot (5) 0.531 0.205 0.449 0.244 0.167 0.271 0.055 0.009 0.202 0.000
@@
Failure Rate GPT 4 Turbo Claude 2 GPT 3.5 Turbo Claude Instant PaLM 2 Yi-34b Mistra-7b MPT-30b Vicuna-13b Phi-1.5
@@
Zeroshot 0.000 0.260 0.000 0.400 0.110 0.330 0.070 1.000 0.480 1.000
Fewshot (-5) 0.000 0.060 0.020 0.000 0.000 0.940 0.480 0.160 0.640 0.860
Fewshot (-4) 0.000 0.033 0.000 0.000 0.017 0.917 0.617 0.117 0.633 0.900
Fewshot (-3) 0.000 0.057 0.057 0.014 0.000 0.871 0.471 0.043 0.600 0.914
Fewshot (-2) 0.000 0.075 0.025 0.000 0.013 0.888 0.538 0.063 0.613 0.888
Fewshot (-1) 0.000 0.044 0.022 0.000 0.011 0.911 0.622 0.078 0.589 0.933
Fewshot (0) 0.000 0.060 0.020 0.000 0.020 0.300 0.380 0.040 0.540 0.880
Fewshot (1) 0.000 0.040 0.020 0.010 0.010 0.290 0.420 0.060 0.520 0.830
Fewshot (2) 0.000 0.040 0.040 0.000 0.010 0.290 0.440 0.030 0.640 0.910
Fewshot (3) 0.000 0.020 0.050 0.010 0.030 0.280 0.460 0.030 0.650 0.830
Fewshot (4) 0.000 0.050 0.030 0.000 0.040 0.280 0.530 0.080 0.630 0.850
Fewshot (5) 0.000 0.050 0.040 0.000 0.020 0.680 0.440 0.050 0.650 0.900

### 12 Jan 2024 v3

@@ Abstract
Complex reasoning ability is one of the most important features of current Large Language Models (LLMs), which has also been leveraged to play an integral role in complex decision-making tasks. Therefore, the investigation into the reasoning capabilities of LLMs is critical: numerous benchmarks have been established to assess the reasoning abilities of LLMs. However, current benchmarks are inadequate in offering a rigorous evaluation of the full extent of reasoning abilities that LLMs are capable of achieving. They are also prone to the risk of overfitting, as these benchmarks, being publicly accessible and static, allow models to potentially tailor their responses to specific benchmark metrics, thereby inflating their performance. Addressing these limitations, our research introduces a new benchmark, named NPHardEval. This benchmark is designed to evaluate the reasoning abilities of LLMs across a broad spectrum of 900 algorithmic questions, extending up to the NP-Hard complexity class. These questions are meticulously chosen to represent a wide range of complexity class below the NP-hard complexity class, offering a rigorous measure of the reasoning ability of LLMs. Through this study, we shed light on the current state of reasoning in LLMs, providing an objective and rigorous perspective through the comparison of LLMs’ performance across complex classes. Our findings contribute significantly to understanding the current capabilities of LLMs in reasoning tasks and lay the groundwork for future advancements in enhancing the reasoning abilities of these models. Moreover, this benchmark is designed with a dynamic update mechanism, where the datapoints are refreshed on a monthly basis. Such regular updates play a crucial role in mitigating the risk of LLMs overfitting to the benchmark, promoting a more accurate and reliable assessment of their reasoning capabilities. The benchmark dataset and code of NPHardEval are available at https://github.com/casmlab/NPHardEval.
@@ 1 Introduction
    Reasoning Ability Gap among Foundation Models: How far away are current open-sourced models compared with closed-sourced models? We aim to quantitatively analyze the differences in performance between closed-source models (including as GPT 4 Turbo, Claude 2, GPT 3.5 Turbo, Claude Instant, and PaLM 2) and open-source models (including Vicuna-13b, Yi-34b, Mistral-7b, MPT-30b, and Phi-1.5). This comparison will shed light on the relative strengths and weaknesses of these models in complex reasoning scenarios.
@@
    Robustness of Benchmark Assessments: This study examines whether the frequent updating of algorithmic benchmarks can effectively prevent the risk of “hacking” the benchmark. The dynamic updating of benchmarks is proposed as a strategy to reduce the likelihood of LLMs overfitting to these benchmarks. However, a pertinent question arises: does finetuning LLMs on benchmarks from the previous month lead to overfitting specific problem types? To explore this, we conducted an experiment where three open-source models -— Phi-2, Mistral-7b, and Qwen-14b —- were finetuned on five distinct versions of the benchmarks. The performance of these models was evaluated on two versions of the benchmark, each differing in difficulty level. This approach allowed us to assess whether finetuning enables models to “hack” benchmarks of varying complexity.
  4. 4.
@@ 2.1 Reasoning ability of LLMs
LLMs brown2020language ; chowdhery2023palm ; chung2022scaling have made significant advancements in natural language processing and related fields. Recent research underscores the unprecedented reasoning abilities of LLMs in various fields, from biomedical and human-computer interaction research to humanities and social studies huang2022towards ; hua2023war ; fan2023datachat ; gao2023examining ; li2023hot . It has been discussed that these models exhibit “emergent” behaviors, including the ability to “reason” when they are large enough wei2022emergent ; schaeffer2023emergent . By providing the models with the chain of thoughts with a simple prompt “Let’s think step by step”, these models are able to answer questions with explicit reasoning steps wei2022chain . This has sparked considerable interest in the community since reasoning ability is a hallmark of human intelligence. Various variations of chain-of-thought have been developed to prompt models’ reasoning abilitykojima2022large ; wang2022iteratively ; hua2022system , such as tree of thoughtyao2023tree , graph of thoughtbesta2023graph , self-inspring techniquewang2023recmind .
@@ 2.2 Benchmarks of LLMs’ Performance
The advancement of LLMs has catalyzed the evolution of a range of general-purpose AI technologies, underscoring the importance of accurately assessing these models’ reasoning capabilities. Existing evaluation approaches predominantly rely on datasets comprising human-generated questions and their standard answers. For instance, MMLU hendrycks2020measuring and GAOKAO zhang2023evaluating both utilize human exam questions in their automated evaluations. Additionally, datasets such as the French National Math Exam, Hungarian National High School Exam [^1] https://huggingface.co/datasets/keirp/hungarian_national_hs_finals_exam, and GHOST (Graduate-Level High-Order Skill Tests)frieder2023mathematical are utilized to assess LLMs’ reasoning proficiency. zhu2023dyval proposes a dynamic graph-based reasoning benchmark. These sources aim to ensure the absence of data leakage. Nonetheless, the requirement for manual verification of answers in these datasets limits their practical utility. In general, these datasets are commonly employed as benchmarks in the field; however, they lack a quantitative metric for assessing the difficulty level of the questions and the extent of reasoning necessary to answer them. This absence of precise measurement criteria results in a limited understanding of the logical reasoning capabilities of large language models.
@@ 3.2 Difficulty Level for Tasks
NPHardEval categorizes the challenges it presents into a hierarchy of difficulty, spanning from the simplest to the most complex. This structure is divided into 10 distinct levels of difficulty for each task, with the initial level being designed as the most basic challenge that an LLM might face. This gradation allows for a nuanced assessment of an LLM’s problem-solving abilities across a spectrum of increasingly complex tasks. For instance, the GCP-D problem has difficulty levels 1 to 10 with questions of 6, 8, 10, 12, 14, 16, 18, 20, 22, and 24 average edges and 6, 7, 8, 9, 10, 11, 12, 13, 14, and 15 nodes. Beginning with graphs of 6 nodes and 6 edges, each subsequent level incorporates an additional 2 edges and 1 node, culminating in graphs of 24 edges and 15 nodes at the most challenging level.
@@ 4 Experimental Setting
This section presents the experiment setting to answer the three research questions. Our approach assessed 10 distinct LLMs, with a dichotomy between five proprietary (closed-source) models and five open-source models, including GPT 4 Turbo, Claude 2, GPT 3.5 Turbo, Claude Instant 1.2, PaLM 2, Vicuna-13b, Yi-34b, Mistral-7b, MPT-30b, and Phi-1.5.
@@ 4.2 Experiment 2: Model Performance across Task Complexity
4.3 Experiment 3: Sanity Check of Finetuning’s Effect on Benchmark Robustness
The primary objective of this experiment is to ascertain whether it is possible to “hack” our benchmark by finetuning models on its previous versions. To simulate this, we constructed five versions of the benchmark, maintaining a consistent difficulty level. Additionally, we utilize two distinct versions of the benchmark, each varying in difficulty, to evaluate the potential for hacking under varying conditions. To replicate the progression of time, models were finetuned sequentially on one to five benchmarks, each finetuned checkpoint is tested on the three distinct benchmarks for evaluation.
The experiment involved finetuning three high-performing open-source models: Phi-2, Mistral-7b, and Qwen-14b. Due to constraints in computing resources, the Yi-34b model was not included in the finetuning process. For the finetuning process, we employed the QLoRA technique, applying specific hyperparameters: batch size set to 8, a single epoch, a warmup proportion of 0.03, a learning rate of 1e-4, lora_r at 64, lora_alpha at 16, and a lora_dropout of 0.1. This approach aims to rigorously test the robustness of our benchmark against potential overfitting strategies.
4.4 Experiment 4: Comparative Analysis of Learnability by In-context Learning
@@ 5.1 Reasoning Ability of Foundation Models
Experiment 1 focuses on a comprehensive comparison among various foundation models. In Figure 2, we present the overall zero-shot accuracy for each problem, providing a visual representation of the performance of different models. Our observations reveal that closed-source models generally demonstrate higher accuracy and a lower rate of failure compared to their open-source counterparts. Notably, GPT-4 Turbo often emerges as the frontrunner in performance across the majority of tasks, indicating its superior problem-solving capabilities, while Claude 2, on the other hand, often performs the best on medium-level (NP-complete) complexity in zero-shot settings. Within the realm of open-source models, Yi-34b, Qwen-14b, and Mistral-7b distinguish themselves by significantly outperforming other models in this category. This disparity highlights a notable performance gap and suggests that these two models possess more advanced reasoning abilities compared to other open-source options.
@@
As Figure 3(a) shows, upon analysis of the weighted accuracy for different models across problem complexities, we observed a general trend where all models experienced a decrease in accuracy as problem complexity increased. Notably, there are two detailed findings for overall reasoning ability change. First, regarding the performance decay speed, among the 12 models we tested, the average performance demonstrated a higher accuracy at the P and NP-Complete complexity levels (with similar weighted accuracies of 0.24 and 0.25) but saw a sharper decline as the problems became more complex when proceeding to the NP-hard level (with a weighted accuracy of 0.02). There is a performance decay on average when models are tested against NP-Hard problems. Second, close-source models usually perform better than open-source models – there are more triangles in the upper locations than squares in Figure 3(a).
Figure 3: Model performance on different complexity problems: (a) weighted accuracy (b) (weighted) failure rate. Open models are denoted in squares and close models are denoted in triangles. Trends of metrics are demonstrated for models with outstanding performances in both weighted accuracy and failure rate, including both close-source (GPT 4 Turbo and Claude 2) and open-source (Mistral-7B and Phi-2) models.
@@ 5.2 Reasoning Performance across Task Complexity and Difficulty Levels
Figure 4: Model’s performance on each complexity level. (a) GPT 4 Turbo. (b) Claude 2. (c) GPT 3.5 Turbo. (d) Claude Instant 1.2. (e) PaLM 2. (f) Yi-34b. (g) Qwen-14b. (h) Mistral-7b. (i) Phi-2. (j) MPT-30b. (k) Vicuna-13b. (l) Phi-1.5.
Figure 5: Model’s performance on tasks across complexity levels. (a) GPT 4 Turbo. (b) Claude 2. (c) GPT 3.5 Turbo. (d) Claude Instant 1.2. (e) PaLM 2. (f) Yi-34b. (g) Qwen-14b. (h) Mistral-7b. (i) Phi-2. (j) MPT-30b. (k) Vicuna-13b. (l) Phi-1.5.
Figure 6 presents the accuracy of each model across various problems associated with P, NP-Complete, and NP-Hard complexities. Regarding P complexity, notable differences emerged among the models. GPT 3.5 Turbo, GPT 4 Turbo, Yi-34b, and Qwen-14b models exhibited significantly superior performance on the SAS problem compared to the other two problems. GPT 3.5 Turbo, Yi-34b, and Vicuna-13b models demonstrated markedly better performance on the EDP problem compared to the SPP problem. Only the Vicuna-13b model displayed slightly better performance, although not significant, on the EDP problem compared to SAS across all investigated models.
Other observations include: GPT 4 Turbo showcased very similar performance between the EDP and SPP problems, while Claude Instant 1.2 exhibited similar performance for all these three problems. Yi-34b, Owen-14b, GPT 3.5 Turbo, and GPT 4 Turbo displayed remarkably high accuracy specifically for the SAS task. MPT-30b and Phi-1.5 showed very limited performance in identifying these three problems.
Regarding NP-Complete complexity, there are several observations to highlight. Still, neither MPT-30b nor Phi-1.5 could deliver any identification for problems in the NP-Complete complexity. In the case of GCP-D and TSP-D problems, the performance of these models varied significantly. Phi-2, Vicuna-13b and GPT 4 Turbo outperformed in the GCP-D problem compared to TSP-D, whereas Claude Instant 1.2, Claude 2, and PaLM 2 exhibited better performance in TSP-D over GCP-D. On the other hand, models like Mistral-7b, Yi-34b, Qwen-14b, and GPT 3.5 Turbo showcased relatively similar performance between these two tasks. For the KSP task, only GPT 4 Turbo demonstrated promising performance, while the remaining models faltered.
@@
5.3 Evaluating Benchmark Vulnerability to Finetuning Strategies
Figure 6: Model’s robustness on different problems and difficulty levels.
In this study, we explore the robustness of benchmark against hacking attempts through a process of finetuning on pairs of question and gold answer. We experiment using 3 well-performing open-source models: Qwen-14b, Mistral-7b, and Phi-2 on two versions of benchmarks. Specifically, we present their performance on two algorithmic questions: the polynomial-time solvable SPP problem and the NP-complete KSP Problem.
Figure 6 presents four graphs illustrating these scenarios: (1) the upper left graph shows SPP challenges at difficulty levels 1-10; (2) the lower left graph displays SPP at levels 11-20; (3) the upper right graph represents KSP at levels 1-10; and (4) the lower right graph shows KSP at levels 11-20. The first row of indicate the accuracy mean of each model, which is averaged over outcomes at 5 finetuning checkpoints, ranging from zero (no finetuning) to five, all within difficulty levels 1-10.
Our findings are twofold: (1) While finetuning yields improvements in solving SPP problems, its impact on the more complex NP-complete KSP is minimal. This suggests the inherent difficulty of hacking NP-complete, and potentially NP-hard, problems through basic question-and-answer approaches. Manual annotation of the thought process, which is not provided in the benchmarks, could potentially enhance effectiveness, albeit with challenges in annotation. (2) Finetuning appears beneficial for performance within the same difficulty level for P problems, yet shows limited out-of-distribution (OOD) adaptability and struggles to generalize to more challenging problems (as evidenced in graphs a and c). For instance, Qwen-14b demonstrates notable proficiency on SPP challenges at levels 1-10 following finetuning; its performance is comparable to that of GPT-4. However, its efficacy significantly diminishes on SPP problems at levels 11-20, even underperforming compared to its unfinetuned state. This indicates that finetuning on these benchmarks could actually impede generalization capabilities and increase the complexity of hacking.
In conclusion, our benchmarks are challenging to hack due to two primary factors: (1) the inherent complexity of NP-complete and NP-hard problems, which are difficult to learn solely from question-and-answer pairs, and (2) the propensity for P problems to become overfitted through finetuning on these pairs, which can be easily exposed when increasing the problem difficulty. Based on these conclusions, we intend to periodically update our benchmarks strategetically with varying difficulty levels to minimize the potential for hacking.
@@ 5.4 Effects of Few-shot Examples’ Difficulty on Reasoning Ability Enhancement
In Experiment 4, we focused on the tasks of SAS and EDP to investigate the nature of the in-context learning capabilities of LLMs. This experiment empirically distinguishes between “learning” and “mimicking” as exhibited by LLMs during in-context learning scenarios. Our findings also revealed a clear dichotomy in the approach to learning and generalization from examples between closed-source and open-source models.
@@
Accuracy GPT 4 Turbo Claude 2 GPT 3.5 Turbo Claude Instant PaLM 2 Yi-34b Qwen-14b Mistral-7b Phi-2 MPT-30b Vicuna-13b Phi-1.5
@@
Zeroshot 1.000 0.445 0.942 0.442 0.416 0.620 0.706 0.149 0.191 0.000 0.113 0.000
Fewshot (-5) 0.978 0.685 0.920 0.735 0.603 0.065 0.103 0.043 0.095 0.165 0.085 0.155
Fewshot (-4) 1.000 0.662 0.902 0.667 0.516 0.093 0.189 0.129 0.149 0.156 0.084 0.118
Fewshot (-3) 0.982 0.694 0.831 0.769 0.496 0.143 0.114 0.153 0.116 0.131 0.067 0.084
Fewshot (-2) 1.000 0.771 0.910 0.710 0.617 0.102 0.070 0.094 0.073 0.117 0.037 0.087
Fewshot (-1) 0.987 0.770 0.896 0.589 0.607 0.048 0.126 0.085 0.107 0.157 0.087 0.057
Fewshot (0) 0.984 0.671 0.846 0.651 0.660 0.598 0.255 0.413 0.222 0.258 0.089 0.098
Fewshot (1) 0.991 0.580 0.878 0.696 0.455 0.593 0.455 0.386 0.287 0.233 0.055 0.109
Fewshot (2) 1.000 0.675 0.829 0.587 0.656 0.647 0.444 0.296 0.260 0.175 0.067 0.056
Fewshot (3) 0.993 0.736 0.800 0.598 0.489 0.662 0.427 0.318 0.275 0.144 0.093 0.098
Fewshot (4) 1.000 0.729 0.869 0.580 0.471 0.638 0.251 0.287 0.195 0.269 0.053 0.106
Fewshot (5) 1.000 0.671 0.844 0.602 0.607 0.167 0.387 0.356 0.196 0.202 0.055 0.064
@@
Zeroshot 0.536 0.120 0.318 0.176 0.033 0.166 0.269 0.058 0.009 0.002 0.147 0.000
Fewshot (-5) 0.387 0.075 0.417 0.048 0.170 0.000 0.000 0.015 0.210 0.000 0.000 0.205
Fewshot (-4) 0.556 0.209 0.367 0.102 0.207 0.000 0.000 0.000 0.300 0.000 0.044 0.284
Fewshot (-3) 0.500 0.178 0.386 0.167 0.235 0.029 0.000 0.029 0.327 0.000 0.108 0.331
Fewshot (-2) 0.462 0.173 0.479 0.210 0.208 0.146 0.065 0.090 0.329 0.000 0.154 0.335
Fewshot (-1) 0.485 0.200 0.513 0.246 0.289 0.135 0.069 0.098 0.348 0.011 0.248 0.328
Fewshot (0) 0.518 0.209 0.564 0.253 0.238 0.282 0.227 0.182 0.320 0.022 0.164 0.293
Fewshot (1) 0.535 0.184 0.535 0.355 0.205 0.089 0.266 0.089 0.115 0.013 0.160 0.115
Fewshot (2) 0.545 0.209 0.544 0.238 0.196 0.195 0.266 0.042 0.098 0.015 0.093 0.087
Fewshot (3) 0.536 0.189 0.449 0.315 0.182 0.127 0.140 0.067 0.060 0.007 0.191 0.051
Fewshot (4) 0.538 0.209 0.507 0.305 0.200 0.247 0.186 0.095 0.009 0.000 0.129 0.000
Fewshot (5) 0.531 0.205 0.449 0.244 0.167 0.271 0.146 0.055 0.015 0.009 0.202 0.000
@@
Failure Rate GPT 4 Turbo Claude 2 GPT 3.5 Turbo Claude Instant PaLM 2 Yi-34b Qwen-14b Mistral-7b Phi-2 MPT-30b Vicuna-13b Phi-1.5
@@
Zeroshot 0.000 0.260 0.000 0.400 0.110 0.330 0.200 0.070 0.150 1.000 0.480 1.000
Fewshot (-5) 0.000 0.060 0.020 0.000 0.000 0.940 0.900 0.480 0.920 0.160 0.640 0.860
Fewshot (-4) 0.000 0.033 0.000 0.000 0.017 0.917 0.817 0.617 0.867 0.117 0.633 0.900
Fewshot (-3) 0.000 0.057 0.057 0.014 0.000 0.871 0.886 0.471 0.886 0.043 0.600 0.914
Fewshot (-2) 0.000 0.075 0.025 0.000 0.013 0.888 0.913 0.538 0.900 0.063 0.613 0.888
Fewshot (-1) 0.000 0.044 0.022 0.000 0.011 0.911 0.856 0.622 0.878 0.078 0.589 0.933
Fewshot (0) 0.000 0.060 0.020 0.000 0.020 0.300 0.640 0.380 0.670 0.040 0.540 0.880
Fewshot (1) 0.000 0.040 0.020 0.010 0.010 0.290 0.500 0.420 0.700 0.060 0.520 0.830
Fewshot (2) 0.000 0.040 0.040 0.000 0.010 0.290 0.510 0.440 0.710 0.030 0.640 0.910
Fewshot (3) 0.000 0.020 0.050 0.010 0.030 0.280 0.450 0.460 0.670 0.030 0.650 0.830
Fewshot (4) 0.000 0.050 0.030 0.000 0.040 0.280 0.570 0.530 0.700 0.080 0.630 0.850
Fewshot (5) 0.000 0.050 0.040 0.000 0.020 0.680 0.520 0.440 0.740 0.050 0.650 0.900